# PRISM: Proof-Carrying Artifact Generation through LLM $\times$ MDE Synergy and Stratified Constraints

### v1

@@ Abstract.
PRISM unifies Large Language Models with Model-Driven Engineering to generate regulator-ready artifacts and machine-checkable evidence for safety- and compliance-critical domains.
PRISM integrates three pillars: a Unified Meta-Model (UMM) reconciles heterogeneous schemas and regulatory text into a single semantic space; an Integrated Constraint Model (ICM) compiles structural and semantic requirements into enforcement artifacts including generation-time automata (GBNF, DFA) and post-generation validators (e.g., SHACL, SMT); and Constraint-Guided Verifiable Generation (CVG) applies these through two-layer enforcement—structural constraints drive prefix-safe decoding while semantic/logical validation produces machine-checkable certificates. When violations occur, PRISM performs audit-guided repair and records generation traces for compliance review.
We evaluate PRISM in automotive software engineering (AUTOSAR) and cross-border legal jurisdiction (Brussels I bis). PRISM produces structurally valid, auditable artifacts that integrate with existing tooling and substantially reduce manual remediation effort, providing a practical path toward automated artifact generation with built-in assurance.
@@ 1.1. From PIM–PSM Transformations to Unified Meta-Models
Model-driven architecture (MDA) advocated raising design effort from code to high-level platform-independent models (PIMs), then systematically refining them into platform-specific models (PSMs) and code (Object Management Group, 2003; Brambilla et al., 2017; Mardani Korani et al., 2023). Recent advances in meta-model integration aim to provide unified views across heterogeneous domain descriptions by merging multiple domain-specific languages (DSLs) into a single unified meta-model (UMM) (Amaral et al., 2023). However, substantial human curation is still required to align concepts and avoid semantic drift during integration, making the “unification power of models” an active research challenge rather than a solved problem.
@@ 1.2. Conformance and Assurance in Safety-Critical Domains
Safety-critical domains demand rigorous conformance to normative standards and bidirectional traceability across the entire development lifecycle (de la Vara et al., 2016). In automotive and avionics practice, models such as AUTOSAR (AUTomotive Open System ARchitecture) XML (ARXML) or the Architecture Analysis & Design Language (AADL) must satisfy prescribed structural and semantic rules; certification bodies additionally require that every requirement, design element, code unit, and test case be traceable in both directions (de la Vara et al., 2016).
Despite mature certification processes, a persistent verification gap remains between model-level properties and generated artifacts: most industrial generators still do not emit machine-checkable correctness certificates. Formal verification—including techniques encouraged for Design Assurance Level (DAL) A/B avionics software via DO-333—is increasingly applied to prove the absence of runtime errors or to discharge temporal-logic obligations. Qualified code generators such as Simulink/Stateflow and SCADE seek to preserve verified properties when emitting production code (Leroy, 2009). Together, these practices accumulate objective evidence for certification authorities, but they rely on a toolchain that is expensive to qualify and difficult to evolve.
@@ 1.3. Pain Points of Conventional MDE Pipelines
Industrial model-driven engineering (MDE) pipelines expose three long-standing limitations.
(i) Evolution overhead. Large, chained transformations are brittle: schema updates or new domain rules induce schema drift, invalidating hard-coded mappings and templates and forcing costly rework.
(ii) Lack of correctness-by-construction. Unlike verified research compilers such as CompCert (Leroy, 2009), which deliver machine-checkable proofs alongside the generated object code, most model compilers emit artifacts with no accompanying certificate. As a result, correctness must be established post hoc through manual review or downstream analysis (Mugnier et al., 2025).
(iii) Missing solver-based self-repair. When a generated artifact violates a constraint, engineers typically diagnose and repair the issue by hand. Automated, constraint-guided repair that spans structural, semantic, and logical layers remains largely confined to research prototypes (Mündler et al., 2025; HouXinyi et al., 2024). These gaps hinder continuous compliance and force recurring review cycles to re-establish conformance after each change.
@@ 1.4. LLMs and Knowledge-Guided Generation: A New Opportunity
Large Language Models (LLMs) such as Codex and GPT, together with open-source peers including StarCoder and DeepSeek-Coder, can already generate compilable multi-file projects directly from natural-language prompts, albeit with varying degrees of structural and semantic correctness (Liu et al., 2024; Idrisov and Schlippe, 2024; Wong et al., 2023). Recent work has begun exploring the synergy between LLMs and Model-Driven Engineering: Lebioda et al. (Petrovic et al., 2024) demonstrate LLM-assisted automotive software development through Ecore model instance creation followed by OCL constraint validation, while Alaoui Mdaghri et al. (Alaoui Mdaghri et al., 2025) propose leveraging LLMs for DSL modeling from natural language descriptions within an iterative validation framework. Similarly, Patil et al. (Patil et al., 2025) show that specification-driven LLM code generation combined with formal verification can produce safety-critical embedded automotive software even without iterative backprompting (Patil et al., 2025). However, unconstrained sampling frequently violates syntax rules or domain-specific constraints.
Grammar- and prefix-constrained decoding mitigates these failures by intersecting the model’s token distribution with a formal grammar, ensuring that every incremental prefix remains well formed while incurring negligible loss in fluency. This has been demonstrated by incremental and semantically constrained approaches such as PICARD and Synchromesh (Scholak et al., 2021; Poesia et al., 2021), and further optimized for subword alignment and efficiency by DOMINO (Beurer-Kellner et al., 2024; Geng et al., 2024).
Beyond syntax enforcement, knowledge-graph-augmented prompting injects authoritative domain triples—for example, AUTOSAR component hierarchies—into the LLM context, grounding generation and reducing hallucinations (Zhang et al., 2024). Finally, iterative LLM+SMT loops are emerging: an LLM proposes an artifact, an SMT solver checks constraints, counterexamples are fed back, and the model (or an auxiliary synthesizer) repairs defects until the artifact satisfies all checks (Tang et al., 2024). Early studies report dramatic reductions in constraint violations, pointing toward pipelines in which artifacts are generated under explicit, checkable constraints—a capability largely missing from classical MDE practice.
@@ 2.1. Model Transformations in MDA
Model-Driven Architecture (MDA) employs model-to-model (M2M) transformations (e.g., ATL, QVT, Triple Graph Grammars (TGG)) and model-to-text (M2T) transformations (e.g., Acceleo, Xtend) to automate structured artifact derivation (Ma et al., 2025a; Kulkarni et al., 2023; Kahani et al., 2019). These transformation languages provide explicit, systematic mappings between abstraction levels, in contrast to ad hoc, manually curated integration efforts that are vulnerable to semantic drift (see Section 1). However, while they improve repeatability, they typically do not offer the kind of explicit semantic preservation guarantees that a colimit-based unified meta-model construction can provide.
@@ 2.2. Specification and Certification in Regulated Domains
Safety-critical sectors require that every development artifact conform to normative schemas and be accompanied by auditable evidence. Standards such as DO-178C and ISO 26262 mandate rigorous conformance and full traceability, and domain-specific coding and safety guidelines—including MISRA-C (Motor Industry Software Reliability Association C guidelines) and SOTIF (Safety Of The Intended Functionality)—require structured assurance cases. Tool qualification (e.g., qualification of Simulink auto-coders), bidirectional trace links, and safety arguments structured using Goal Structuring Notation (GSN) are routinely required; empirical studies document the effort needed to curate evidence that ties each safety claim to verified artifacts (El-Gnainy et al., 2024; Nair et al., 2014).
As a result, model-driven toolchains in automotive, avionics, and healthcare often embed schema validators and certification artifact generators to ensure that outputs such as AUTOSAR XML and AADL designs satisfy domain rules “by construction.” This reduces the manual burden at audit time but further entrenches specialized, domain-qualified tooling that is costly to evolve.
@@ 2.3. Formal Verification and Proof-Carrying Artifacts
Formal methods enrich modeling languages with machine-checkable semantics. Contracts in the form of assume/guarantee pairs and Object Constraint Language (OCL) invariants enable compositional reasoning; model checkers such as SPIN and NuSMV, as well as Satisfiability Modulo Theories (SMT)-based analyzers, are used to validate behavioral properties before code generation. Interactive proof assistants (e.g., Isabelle, Coq, PVS) remain essential wherever deep functional guarantees are required.
The proof-carrying code (PCC) paradigm attaches machine-verifiable correctness certificates to generated artifacts, providing a theoretical foundation for safety-critical practice, though complete PCC-style adoption in industrial pipelines is still limited. Solver-aided synthesis and repair techniques push this further: Max-SAT-based approaches such as DirectFix compute minimal patches that re-establish property satisfaction (Orvalho et al., 2024; Mechtaev et al., 2015).
Recent domain studies also demonstrate scalability (Charalambous et al., 2023). For instance, AUTOSAR XML models have been translated into timed-automata templates and exhaustively analyzed with UPPAAL for end-to-end latency verification (Zhang et al., 2023), and Architecture Analysis & Design Language (AADL) models enriched with a Safety Annex have undergone model-checking-driven fault injection to enumerate minimal cut sets (Stewart et al., 2021). The PRISM pipeline—which integrates the Unified Meta-Model (UMM), Integrated Constraint Model (ICM), and Constraint-Guided Verifiable Generation (CVG) stages—inherits this “lift-to-formal-semantics” philosophy while additionally attaching proof-carrying certificates to each generated artifact (Ma et al., 2025b).
@@ 2.4. LLM-Assisted Generation under Structural Constraints
To address this, prior work explores grammar- and prefix-constrained decoding, in which decoding proceeds under a context-free grammar or a deterministic automaton so that every incremental prefix remains syntactically valid, as demonstrated by Synchromesh and PICARD (Scholak et al., 2021), with negligible fluency loss. An alternative strategy, grammar prompting (Wang et al., 2023b), guides the LLM to first synthesize a minimally sufficient grammar tailored to the specific generation task, and then to generate content under that task-specific grammar. This treats constraint specification as an LLM-interpretable guide rather than a hard automaton mask.
@@
More recent decoding algorithms such as DOMINO enforce constraints at the subword level and combine speculative decoding with precomputation to achieve near-zero runtime overhead while still honoring a target grammar (Beurer-Kellner et al., 2024). Follow-up work on fast grammar-constrained decoding shows how to align grammar tokens with the model tokenizer and to build the automaton efficiently, reducing offline preprocessing cost by more than an order of magnitude while preserving efficient online masking (Park et al., 2025b).
@@
In parallel, reinforcement-learning-style schema alignment aims to train LLMs to emit JSON that is provably valid with respect to large, real-world schemas. Nevertheless, even frontier models systematically fail on complex industrial schemas, which has motivated dedicated benchmarks such as SchemaBench and analyses such as JSONSchemaBench that characterize the trade-off between “over-constrained” engines (which reject semantically valid but grammatically atypical structures) and “under-constrained” engines (which accept ill-formed outputs and thus require repair) (Lu et al., 2025). Retrieval- and knowledge-graph-augmented prompting grounds generation in authoritative domain ontologies (for example, AUTOSAR component hierarchies), mitigating hallucinations. Finally, Ferrari and Spoletini show that LLMs can synthesize first-order temporal predicates and safety conditions directly from regulatory prose, effectively extracting normative logical constraints from natural-language standards (Ferrari and Spoletini, 2025).
@@
Emerging industrial pipelines iterate on these trends: an LLM drafts an artifact; for structural layers, a supported subset of JSON Schema / regular expressions / Generalized Backus–Naur Form (GBNF) is compiled into deterministic automata to guide decoding via prefix-safe masking and online tracing; post-generation semantic and logic validators (e.g., SHACL, SMT) then check semantic and logical constraints, capture machine-checkable evidence, and drive automated repair. PRISM generalizes this pattern into a unified, auditable “verify-as-you-generate” loop.
@@
A critical challenge identified in constrained decoding research is distribution distortion: while grammar masking guarantees structural validity, it can bias LLM generation toward the shortest valid completions, yielding artifacts that are syntactically correct but semantically impoverished (Park et al., 2025a). This effect arises because hard constraint masking eliminates longer yet equally valid continuations, thereby distorting the LLM’s learned probability distribution.
@@
Grammar-Aligned Decoding (GAD) (Park et al., 2025a) addresses this through adaptive sampling with approximate expected futures (ASAp): rather than deterministically masking all invalid tokens, GAD performs limited forward simulation to estimate which candidate tokens can still lead to a grammatical completion, and samples from a distribution that balances grammaticality with fidelity to the base LLM’s preferences. Empirical results on code generation and structured data extraction show that GAD produces outputs with higher likelihood under the base model while maintaining grammatical correctness.
@@ 3.1. Overview
Figure 1. PRISM links domain requirements to verifiable generation. Natural-language requirements and existing meta-models are consolidated into a Unified Meta-Model (UMM) and an Integrated Constraint Model (ICM). Generation is guided by structural constraints (Layer-1, enforced during decoding) and then checked by semantic / logical validators (Layer-2) after decoding. The validation signals, together with the decode-time audit record, form evidence that can be used to automatically repair the artifact in a closed loop.:
@@
Industrial, safety-critical domains (e.g., AUTOSAR-compliant automotive software, legal compliance workflows) require not only functionally plausible artifacts but also verifiably correct artifacts that can be audited by certification authorities (Yang et al., 2024; Wang et al., 2023a). Unconstrained Large Language Models (LLMs) excel at understanding domain intent and producing structured drafts, but they do not by themselves guarantee regulatory conformance, global consistency, or traceable justification. In contrast, Model-Driven Engineering (MDE) offers rigor through meta-models, transformation rules, and formal analysis, yet is brittle to evolving standards and costly to maintain as domains change. PRISM addresses this gap as an end-to-end high-assurance generation pipeline that couples the flexibility of LLMs with the accountability of formal methods. The pipeline is organized around three tightly connected components: (1) the Unified Meta-Model (UMM), which provides an auditable semantic backbone across heterogeneous domain sources; (2) the Integrated Constraint Model (ICM), which aggregates structural, semantic, and logical constraints; and (3) the Constraint-guided Verifiable Generation (CVG) process, which enforces constraints during generation, certifies artifacts after generation, and drives targeted repair when violations are found. Together, these components allow PRISM to automate the 80% of routine artifact construction while still producing machine-checkable evidence for expert review on the remaining 20%.
@@
PRISM coordinates LLMs and formal methods in a relay pattern: the UMM and ICM formalize what “correct” means, and CVG—which includes Audit-Guided Repair (AGR)—uses those definitions to control generation, verify conformance, and close the loop with targeted repairs when inconsistencies are detected. We summarize each component below.
Unified Meta-Model (UMM). The UMM is not a passive schema repository; it is the semantic contract that both constrains generation and explains to auditors what each generated element is supposed to mean. To build the UMM, PRISM merges heterogeneous meta-models and domain ontologies (e.g., AUTOSAR XSD, legacy UML variants) into a coherent, typed graph $M_{\text{UMM}}=(V,E,\mathcal{C}_{\text{struct}})$, where nodes denote domain entities (classes, attributes) and edges record structural relations. This merge is carried out using a mathematically well-defined unification procedure (instantiated in our implementation as a category-theoretic colimit construction (Diskin and Maibaum, 2012)), which ensures that overlapping concepts from different sources are consistently identified and traceable. The point is not that the colimit construction itself is a contribution, but that it lets us claim semantic coherence and auditability: every node in the UMM can be traced back to its source models. For generation, PRISM materializes relevant subgraphs of the UMM as retrieval-augmented context—e.g., JSON-style schema fragments and textual descriptions—which are injected into LLM prompts to give the model a precise structural and conceptual scaffold.
@@
  • Channel 2 (Inductive, LLM-assisted). Many safety or regulatory requirements in practice are not fully captured by XSDs or UML—for example, “an OperationInvokedEvent must reference an existing operation,” or “mode transitions must form a directed acyclic graph.” Channel 2 uses an LLM extractor $E_{\theta}$ to read natural-language specifications (e.g., AUTOSAR Software Component Template PDFs) that have been augmented with inlined UMM definitions (see Figure 3). $E_{\theta}$ proposes candidate L2-only constraints (expressed as semantic or logic assertions). Before these constraints are admitted into the ICM, PRISM links each predicate back to UMM entities via entity linking (Definition 4) and checks their semantic compatibility (Definition 6), then runs consistency check to reject contradictory rules (Algorithm 2).
This dual-channel design allows PRISM to ingest both structured sources (Channel 1) and free-text specifications (Channel 2) without sacrificing rigor: LLM-proposed rules do not bypass formal scrutiny, but instead become auditable constraints that are grounded in—and indexed by—the UMM.
Constraint-guided Verifiable Generation (CVG). CVG is the execution core of PRISM. It governs how artifacts are produced, checked, and—if necessary—repaired. It proceeds in three stages:

### v2

# ATLAS: Artifact Generation Through Layered Constraints and LLM $\times$ MDE Synergy

@@ Abstract.
ATLAS unifies Large Language Models with Model-Driven Engineering to generate regulator-ready artifacts and machine-checkable evidence for safety- and compliance-critical domains.
ATLAS integrates three pillars: a Unified Meta-Model (UMM) reconciles heterogeneous schemas and regulatory text into a single semantic space; an Integrated Constraint Model (ICM) extends our prior Dual-Stage (S2D2) extraction logic to compile layered requirements into deterministic generation-time automata (Layer 1) and post-generation validators (Layer 2); and Constraint-Guided Verifiable Generation (CVG) applies these through two-layer enforcement—Layer 1 structural constraints drive prefix-safe decoding while Layer 2 semantic/logical validation produces machine-checkable certificates. When violations occur, ATLAS performs audit-guided repair and records generation traces for compliance review.
We evaluate ATLAS in automotive software engineering (AUTOSAR) and cross-border legal jurisdiction (Brussels I bis). ATLAS produces structurally valid, auditable artifacts that integrate with existing tooling and substantially reduce manual remediation effort, validating a graduated automation paradigm that automates routine construction while empowering experts to resolve complex semantic ambiguities through machine-checkable evidence.
@@ 1.1. From PIM–PSM Transformations to Unified Meta-Models
Model-driven architecture (MDA) advocated raising design effort from code to high-level platform-independent models (PIMs), then systematically refining them into platform-specific models (PSMs) and code (omg2003mda; 10.5555/3103551; mardanikoraniModelDrivenEngineeringTechniques2023). Recent advances in meta-model integration aim to provide unified views across heterogeneous domain descriptions by merging multiple domain-specific languages (DSLs) into a single unified meta-model (UMM) (amaralUsingMetamodelComposition2023). However, substantial human curation is still required to align concepts and avoid semantic drift during integration, making the “unification power of models” an active research challenge rather than a solved problem.
@@ 1.2. Conformance and Assurance in Safety-Critical Domains
Safety-critical domains demand rigorous conformance to normative standards and bidirectional traceability across the entire development lifecycle (delavaraModelbasedSpecificationSafety2016). In automotive and avionics practice, models such as AUTOSAR (AUTomotive Open System ARchitecture) XML (ARXML) or the Architecture Analysis & Design Language (AADL) must satisfy prescribed structural and semantic rules; certification bodies additionally require that every requirement, design element, code unit, and test case be traceable in both directions (delavaraModelbasedSpecificationSafety2016).
Despite mature certification processes, a practical trade-off remains between assurance and agility: most industrial toolchains rely on rigorous tool qualification rather than per-artifact verification. While effective, this process-heavy approach assumes the generator is static and trusted, creating a toolchain that is expensive to qualify and difficult to evolve when standards or requirements change.
@@ 1.3. Pain Points of Conventional MDE Pipelines
Industrial MDE pipelines suffer from three long-standing limitations: (i) Evolution overhead, where brittle chained transformations require costly rework upon schema updates; (ii) Rigidity of assurance mechanisms, where relying on pre-qualified tools restricts the ability to adopt new features or custom heuristics without breaking the chain of trust, often forcing developers back into manual review for unsupported constraints; and (iii) Missing solver-based self-repair, where automated repair spanning structural and logical layers remains confined to research prototypes (mundlerTypeConstrainedCodeGeneration2025; houxinyiLargeLanguageModels2024). These gaps hinder continuous compliance and force recurring review cycles.
@@ 1.4. LLMs and Knowledge-Guided Generation: A New Opportunity
Large Language Models (LLMs) such as Codex and GPT, together with open-source peers including StarCoder and DeepSeek-Coder, can already generate compilable multi-file projects directly from natural-language prompts, albeit with varying degrees of structural and semantic correctness (liuNoNeedLift2024; Idrisov2024; Wong2023). Recent work has begun exploring the synergy between LLMs and Model-Driven Engineering: Lebioda et al. (Petrovic_2024) demonstrate LLM-assisted automotive software development through Ecore model instance creation followed by OCL constraint validation, while Alaoui Mdaghri et al. (Alaoui2025) propose leveraging LLMs for DSL modeling from natural language descriptions within an iterative validation framework. Similarly, Patil et al. (Patil2025) show that specification-driven LLM code generation combined with formal verification can produce safety-critical embedded automotive software even without iterative backprompting (Patil2025). However, unconstrained sampling frequently violates syntax rules or domain-specific constraints.
Grammar- and prefix-constrained decoding mitigates these failures by intersecting the model’s token distribution with a formal grammar, ensuring that every incremental prefix remains well formed while incurring negligible loss in fluency. This has been demonstrated by incremental and semantically constrained approaches such as PICARD and Synchromesh (scholakPICARDParsingIncrementally2021; poesiaSynchromeshReliableCode2021), and further optimized for subword alignment and efficiency by DOMINO (zotero-item-1597; gengGrammarConstrainedDecodingStructured2024).
Beyond syntax enforcement, knowledge-graph-augmented prompting injects authoritative domain triples—for example, AUTOSAR component hierarchies—into the LLM context, grounding generation and reducing hallucinations (zhangKnowGPTKnowledgeGraph2024). Finally, iterative feedback-driven refinement is emerging: an LLM proposes an artifact, external validators provide diagnostics, and the model repairs defects based on this feedback (tangCodeRepairLLMs2024). This aligns with recent findings on “verbal reinforcement learning” and self-debugging (shinn2023reflexion; chen2024selfdebug), pointing toward pipelines in which artifacts are generated under explicit, checkable constraints. Early studies report dramatic reductions in constraint violations, pointing toward pipelines in which artifacts are generated under explicit, checkable constraints—a capability largely missing from classical MDE practice.
@@ 2.1. Model Transformations in MDA
Model-Driven Architecture (MDA) automates structured artifact derivation via model-to-model and model-to-text transformations (DualStageFramework; kulkarniSymbioticApproachLeveraging2023; kahaniSurveyClassificationModel2019). While these systematic mappings improve repeatability over ad hoc integration (see Section 1), they typically suffer from inherent rigidity and platform dependence, making it difficult to inject custom logic or to cost-effectively construct meta-models for domains lacking explicit schemas. Our previous work, S2D2 (DualStageFramework), addressed some of these rigidities by introducing Dual-Path Fusion for static structure generation. ATLAS evolves this by moving from static code generation to dynamic constraint enforcement for LLMs.
@@ 2.2. Specification and Certification in Regulated Domains
Safety-critical sectors mandate rigorous conformance to standards (e.g., DO-178C, ISO 26262) and structured assurance cases (e.g., MISRA-C, SOTIF). Empirical studies highlight the significant effort required to curate auditable evidence linking safety claims to artifacts (el-gnainyAIEnhancedAUTOSARConfiguration2024a; nairExtendedSystematicLiterature2014). Consequently, industrial toolchains often embed rigid schema validators to ensure “correctness-by-construction,” reducing audit burdens but entrenching specialized, costly-to-evolve tooling.
@@ 2.3. Formal Verification and Proof-Carrying Artifacts
Formal methods enrich modeling with machine-checkable semantics via contracts (OCL), model checkers (SPIN, SMT), and interactive proof assistants. The proof-carrying code (PCC) paradigm extends this by attaching correctness certificates, a concept pushed further by solver-aided repair techniques like Max-SAT (orvalhoCounterexampleGuidedProgram2024; mechtaevDirectFixLookingSimple2015). Recent studies demonstrate scalability in domain-specific verification (charalambousNewEraSoftware2023), such as analyzing AUTOSAR latency with UPPAAL (zhangAutomaticModellingVerification2023) or enumerating cut sets in AADL (stewartAADLBasedSafetyAnalysis2021). The ATLAS pipeline inherits this “lift-to-formal-semantics” philosophy by integrating UMM, ICM, and CVG stages to attach proof-carrying certificates to generated artifacts (maSpecGenAutomatedGeneration2025).
@@ 2.4. LLM-Assisted Generation under Structural Constraints
To address this, prior work explores grammar- and prefix-constrained decoding, in which decoding proceeds under a context-free grammar or a deterministic automaton so that every incremental prefix remains syntactically valid, as demonstrated by Synchromesh and PICARD (scholakPICARDParsingIncrementally2021), with negligible fluency loss. An alternative strategy, grammar prompting (wangGrammarPromptingDomainspecific2023), guides the LLM to first synthesize a minimally sufficient grammar tailored to the specific generation task, and then to generate content under that task-specific grammar. This treats constraint specification as an LLM-interpretable guide rather than a hard automaton mask.
@@
More recent decoding algorithms such as DOMINO enforce constraints at the subword level and combine speculative decoding with precomputation to achieve near-zero runtime overhead while still honoring a target grammar (zotero-item-1597). Follow-up work on fast grammar-constrained decoding shows how to align grammar tokens with the model tokenizer and to build the automaton efficiently, reducing offline preprocessing cost by more than an order of magnitude while preserving efficient online masking (Park2025).
@@
In parallel, reinforcement-learning-style schema alignment aims to train LLMs to emit JSON that is provably valid with respect to large, real-world schemas. Nevertheless, even frontier models systematically fail on complex industrial schemas, which has motivated dedicated benchmarks such as SchemaBench and analyses such as JSONSchemaBench that characterize the trade-off between “over-constrained” engines (which reject semantically valid but grammatically atypical structures) and “under-constrained” engines (which accept ill-formed outputs and thus require repair) (Lu2025b). Retrieval- and knowledge-graph-augmented prompting grounds generation in authoritative domain ontologies (for example, AUTOSAR component hierarchies), mitigating hallucinations. Finally, Ferrari and Spoletini show that LLMs can synthesize first-order temporal predicates and safety conditions directly from regulatory prose, effectively extracting normative logical constraints from natural-language standards (ferrariFormalRequirementsEngineering2025).
@@
Emerging industrial pipelines iterate on these trends: an LLM drafts an artifact; for structural layers, a supported subset of JSON Schema / regular expressions / Generalized Backus–Naur Form (GBNF) is compiled into deterministic automata to guide decoding via prefix-safe masking and online tracing; post-generation semantic and logic validators (e.g., SHACL, SMT) then check semantic and logical constraints, capture machine-checkable evidence, and drive automated repair. ATLAS generalizes this pattern into a unified, auditable “verify-as-you-generate” loop.
@@
A critical challenge identified in constrained decoding research is distribution distortion: while grammar masking guarantees structural validity, it can bias LLM generation toward the shortest valid completions, yielding artifacts that are syntactically correct but semantically impoverished (parkGrammaralignedDecoding2025). This effect arises because hard constraint masking eliminates longer yet equally valid continuations, thereby distorting the LLM’s learned probability distribution.
@@
Grammar-Aligned Decoding (GAD) (parkGrammaralignedDecoding2025) addresses this through adaptive sampling with approximate expected futures (ASAp): rather than deterministically masking all invalid tokens, GAD performs limited forward simulation to estimate which candidate tokens can still lead to a grammatical completion, and samples from a distribution that balances grammaticality with fidelity to the base LLM’s preferences. Empirical results on code generation and structured data extraction show that GAD produces outputs with higher likelihood under the base model while maintaining grammatical correctness.
@@ 3.1. Overview
Figure 1. ATLAS links domain requirements to verifiable generation. Natural-language requirements and existing meta-models are consolidated into a Unified Meta-Model (UMM) and an Integrated Constraint Model (ICM). Generation is guided by structural constraints (Layer-1, enforced during decoding) and then checked by semantic / logical validators (Layer-2) after decoding. The validation signals, together with the decode-time audit record, form evidence that can be used to automatically repair the artifact in a closed loop.
Block diagram of the ATLAS pipeline. Inputs (NL requirements and meta-model specs) are merged into a unified meta-model and constraint model. L1 automaton-guided decoding enforces structural constraints during generation, L2 semantic and logic validators check the final artifact, and the audit plus validation results feed an audit-guided repair loop.
@@
Industrial, safety-critical domains (e.g., AUTOSAR-compliant automotive software, legal compliance workflows) require not only functionally plausible artifacts but also verifiably correct artifacts that can be audited by certification authorities (yangEvaluation2024; wangMachineDeepLearning2023). Unconstrained Large Language Models (LLMs) excel at understanding domain intent and producing structured drafts, but they do not by themselves guarantee regulatory conformance, global consistency, or traceable justification. In contrast, Model-Driven Engineering (MDE) offers rigor through meta-models, transformation rules, and formal analysis, yet is brittle to evolving standards and costly to maintain as domains change. ATLAS addresses this gap as an end-to-end high-assurance generation pipeline that couples the flexibility of LLMs with the accountability of formal methods. The pipeline is organized around three tightly connected components: (1) the Unified Meta-Model (UMM), which provides an auditable semantic backbone across heterogeneous domain sources; (2) the Integrated Constraint Model (ICM), which aggregates structural, semantic, and logical constraints; and (3) the Constraint-guided Verifiable Generation (CVG) process, which enforces constraints during generation, certifies artifacts after generation, and drives targeted repair when violations are found. Together, these components allow ATLAS to automate the 80% of routine artifact construction while still producing machine-checkable evidence for expert review on the remaining 20%.
@@
ATLAS coordinates LLMs and formal methods in a relay pattern: the UMM and ICM formalize what “correct” means, and CVG—which includes Audit-Guided Repair (AGR)—uses those definitions to control generation, verify conformance, and close the loop with targeted repairs when inconsistencies are detected. We summarize each component below.
Unified Meta-Model (UMM). The UMM is not a passive schema repository; it is the semantic contract that both constrains generation and explains to auditors what each generated element is supposed to mean. To build the UMM, ATLAS merges heterogeneous meta-models and domain ontologies (e.g., AUTOSAR XSD, legacy UML variants) into a coherent, typed graph $M_{\text{UMM}}=(V,E,T,\mathcal{C}_{\text{struct}})$, where nodes denote domain entities (classes, attributes) and edges record structural relations. This merge is carried out using a mathematically well-defined unification procedure (instantiated in our implementation as a category-theoretic colimit construction (diskin2012category)), which ensures that overlapping concepts from different sources are consistently identified and traceable. The point is not that the colimit construction itself is a contribution, but that it lets us claim semantic coherence and auditability: every node in the UMM can be traced back to its source models. For generation, ATLAS materializes relevant subgraphs of the UMM as retrieval-augmented context—e.g., JSON-style schema fragments and textual descriptions—which are injected into LLM prompts to give the model a precise structural and conceptual scaffold.
@@
  • Channel 2 (Inductive, LLM-assisted). Many safety or regulatory requirements in practice are not fully captured by XSDs or UML—for example, “an OperationInvokedEvent must reference an existing operation,” or “mode transitions must form a directed acyclic graph.” Channel 2 uses an LLM extractor $E_{\theta}$ to read natural-language specifications (e.g., AUTOSAR Software Component Template PDFs) that have been augmented with inlined UMM definitions (see Figure 3). $E_{\theta}$ proposes candidate L2-only constraints (expressed as semantic or logic assertions). Before these constraints are admitted into the ICM, ATLAS links each predicate back to UMM entities via entity linking (Definition 3) and checks their semantic compatibility (Definition 5), then runs consistency check to reject contradictory rules (Algorithm 2).
This dual-channel design allows ATLAS to ingest both structured sources (Channel 1) and free-text specifications (Channel 2) without sacrificing rigor: LLM-proposed rules do not bypass formal scrutiny, but instead become auditable constraints that are grounded in—and indexed by—the UMM.
Constraint-guided Verifiable Generation (CVG). CVG is the execution core of ATLAS. It governs how artifacts are produced, checked, and—if necessary—repaired. It proceeds in three stages: